# Deep Learning for Traffic Flow Prediction using Cellular Automata-based Model and CNN-LSTM architecture

### v1

@@ Abstract
Recent works have attempted to use deep learning to predict future states of traffic flow, but have met with mixed results. These approaches face two key challenges. First, training deep learning neural networks requires large amounts of training data which are not yet easily available for traffic flow systems. Second, even when data is available, the neural networks require access to historical data that covers most possible traffic flow dynamics to successfully predict future traffic states. Specifically, these deep learning approaches do not fully leverage domain-knowledge about traffic flow dynamics, despite a significant existing knowledge-base. In this work, we propose to solve both issues using a Convolutional Neural Network (CNNs) with Long Short Term Memory (LSTM) deep learning architecture to successfully predict traffic flow, while leveraging a cellular automata-based statistical mechanics model of traffic flow to generate training and test data. Another major contribution of this paper is the insight that training data for a large traffic system can actually be sampled from the simulations of a much smaller traffic system. This is achieved through observing that the normalized energy distribution of the statistical mechanics model is scale invariant, which significantly eases the burden of data generation for large scale traffic systems. The resulting simulations indicate good agreement between the predicted and the true traffic flow dynamics.

### v2

# Energy-Guided Data Sampling for Traffic Prediction with Mini Training Datasets

@@ Abstract
Recent endeavors aimed at forecasting future traffic flow states through deep learning encounter various challenges and yield diverse outcomes. A notable obstacle arises from the substantial data requirements of deep learning models, a resource often scarce in traffic flow systems. Despite the abundance of domain knowledge concerning traffic flow dynamics, prevailing deep learning methodologies frequently fail to fully exploit it. To address these issues, we propose an innovative solution that merges Convolutional Neural Networks (CNNs) with Long Short-Term Memory (LSTM) architecture to enhance the prediction of traffic flow dynamics. A key revelation of our research is the feasibility of sampling training data for large traffic systems from simulations conducted on smaller traffic systems. This insight suggests the potential for referencing a macroscopic-level distribution to inform the sampling of microscopic data. Such sampling is facilitated by the observed scale invariance in the normalized energy distribution of the statistical mechanics model, thereby streamlining the data generation process for large-scale traffic systems. Our simulations demonstrate promising agreement between predicted and actual traffic flow dynamics, underscoring the efficacy of our proposed approach.